a certain field is a rectangle with a perimeter of 934 feet.the length is 179 feet more than the width.find the width of the rectanglar field
Answer by basketball1127(19)

Let x = length of field
y = width of field
Then
x = y + 179
2x + 2y = 934
x + y = 467
Substituting,
(y + 179) + y = 467
2y + 179= 467
2y = 467 - 179
2y = 288
y = 144
width = 144ft